The dean at Brookfield College found that 20% of the female students are majoring in engineering. 57% of students at Brookfield are women. Let A be the event that the student is a women and B the event that the student is an engineering major.
a.) Which of the following is true? P(B/A) =.20 P (A and B) =.20
b.) What is the probability that a student is chosen at random will be a women engendering major?
c.) If 27% of Brookfield's students are majoring in engineering, what percentage of engineering majors are women?

Answer by stanbon(75887)

The dean at Brookfield College found that 20% of the female students are majoring in engineering. 57% of students at Brookfield are women. Let A be the event that the student is a women and B the event that the student is an engineering major.
a.) Which of the following is true? P(B/A) =.20 P (A and B) =.20
Ans: P(B|A) = 0.20
-----------------------------------
b.) What is the probability that a student is chosen at random will be a women engendering major?
Ans: P(A and B) = 0.57*0.20 = 11.4%
------------------------------------
c.) If 27% of Brookfield's students are majoring in engineering, what percentage of engineering majors are women?
Ans: 11.4/27 = 0.422 = 42.2%
--------------
Cheers,
Stan H.
-----------------